Two point charges, 1.5 nC each, are located at points (25 m,0) and (-25 m,0). What is the potential at point (0, 2 m)?
Answer #1

distance of point of interest from first two charges = squareroot(252 + 22) = 25.08 m

Potential energy = kq1/r + kq2/r

= 2kq1/r

= 2*9*109*1.5*10-9 / 25.08 = 1.076 V